Calculate the mass of Na2O that forms if 4,8g O2 reacts completely with sodium


1
Expert's answer
2022-04-25T17:11:04-0400

Solution:

The molar mass of O2 is 32 g/mol

Therefore,

Moles of O2 = (4.8 g O2) × (1 mol O2 / 32 g O2) = 0.15 mol O2


Balanced chemical equation:

4Na + O2 → 2Na2O

According to stoichiometry:

1 mol of O2 produces 2 mol of Na2O

Thus, 0.15 mol of O2 produces:

(0.15 mol O2) × (2 mol Na2O / 1 mol O2) = 0.30 mol Na2O


The molar mass of Na2O is 62 g/mol

Therefore,

Mass of Na2O = (0.30 mol Na2O) × (62 g Na2O / 1 mol Na2O) = 18.6 g Na2O

Mass of Na2O = 18.6 g


Answer: 18.6 grams of Na2O